A Bibliography of Publications of Marc Van Droogenbroeck

Marc Van Droogenbroeck Institut Montefiore B28, I 76.a

Sart Tilman B-4000 Liège, Belgium

Tel: +32 (0)4 366 26 93 (secr. +32 (0)4 366 26 91) FAX: +32 (0)4 366 29 89

E-mail: M.VanDroogenbroeck@ulg.ac.be (Internet)

13 October 2017 Version 1.05

Abstract

This bibliography records publications of Marc Van Droogenbroeck.

Title word cross-reference

Acquisition [Van01a]. algorithm [VV92]. Algorithms [Van02a, VB05]. analogiques [Van02b]. Analyse [Van01b]. anchors [VB05]. application [Van94c]. applications [SS94]. arbitrary [VT96].

based [Van95, VD02, VB05]. binary [Van94b, Van02a]. broadband [VVD96].

characterization [MV92]. codage [Van94c]. Coding [BCM^{+}93, MV92, SMV93]. computation [VT96]. conception [Van01b]. content [Van95]. critical [VVD96].

d’algorithmes [Van94c]. Data [DMV98]. deconvolution [VV92]. decorrelation [CMV92, CMV93]. d’images [Van94c]. d’objet [Van94c].

Edge [MV92]. education [VVD96]. elements [VT96, Van02a]. embedding [VD02]. entropy [VD02]. erosions [VB05]. evaluation [VVD96]. extension [Van94b].

Fast [VT96, VB05].

hierarchical [MV92].

image [CMV92, CMV93, SS94, SMV93, Van95, Van01a]. images [Van01c, Van02a, VD02]. implementation [Van94a]. information [VD02]. Integration [VV92, DMV98]. Internet [Van02c].

label [Van02a].

Mathematical [SS94]. mathématique [Van94c]. methodology [Van95]. Morphological [SMV93, VB05, Van94a, VT96]. morphologie [Van94c]. morphology [SS94]. moyen [Van94c]. multimédia [Van02c]. multiresolution [VV92].

Narrowband [VVD96]. networks [DMV98]. notion [Van94c]. numérique [Van01c]. numériques [Van94c, Van02b].

opening [Van94b]. openings [Van02a, VB05]. operations [Van94a, VT96]. ordinateurs [Van02d].

Polymorphism [Van94b]. Principes [Van02b]. processing [CMV92, CMV93, SS94]. prospective [CMV92, CMV93].

rectangular [Van02a]. regions [Van95].

scheme [VV92]. schemes [MV92]. segmentation [Van95]. selective [VV92]. sketch [SMV93]. spectral [Van95]. structuring [VT96, Van02a]. subband [MV92]. systèmes [Van01b].

technique [VD02]. techniques [CMV92, CMV93]. Technologies [Van02c]. tele [VVD96]. tele-education [VVD96]. télécommunications [Van01b, Van02b, Van02c, Van02d]. Traitement [Van94c, Van01c, Van01a].

utilisant [Van94c].

versus [VVD96].

References

Bouchard:1993:CI

[BCM^{+}93] L. Bouchard, J. Cassas, F. Meyer, C. Oddou, L. Torres, and M. Van Droogenbroeck. Coding of inside.

Deliverable D7 of Race Project Morpheco R2053, December 1993.

Comes:1992:RPD

[CMV92] S. Comes, S. Maes, and M. Van Droogenbroeck. Recent and prospective decorrelation techniques for image processing. In International Symposium on Signals, Systems and Electronics, pages 136– 141. URSI, Paris, September 1992.

Comes:1993:RPD

[CMV93] S. Comes, S. Maes, and M. Van Droogenbroeck. Recent and prospective decorrelation techniques for image processing. Annales des T ’el ’ecommunications, 48(7–8):390–403, juillet–août 1993. CODEN AN-TEAU. ISSN 0003-4347.

DePraetere:1998:DNI

[DMV98] J. De Praetere, P. Maricau, and M. Van Droogenbroeck. Data networks integration. In 37th European Telecommunications Congress. FITCE’98. “Diverging roles in a converging marketplace”, pages 128–131. Instn. British Telecommun. Eng., London, UK, August

1998. URL http://www.ulg.ac.
be/telecom/publi/publications/
mvd/fitce1998.pdf.

Maes:1992:ECH

[MV92] S. Maes and M. Van Droogenbroeck. Edge characterization in hierarchical subband coding schemes. In Fourth International Conference on Image Processing and its Applications: 7–9 April 1992, venue, Maastricht Exhibition and Convention Centre, The Netherlands, volume 354 of Conference publication, pages 598–601. IEE, London, UK, May 1992. ISBN 0-85296-543-5. LCCN TA1632 .I553 1992.

Simon:1993:MIS

[SMV93] B. Simon, B. Macq, and M. Van Droogenbroeck. Morphological image sketch coding. In Workshop on Circuits, Systems and Signal Processing, pages 251–256. ProRISC IEEE Benelux, Houthalen, March 1993.

Serra:1994:MMA

[SS94] Jean Serra and Pierre Soille, editors. Mathematical morphology and its applications to image processing, volume 2 of Computational imaging and vision. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-3093-5. LCCN TA1637 .M37 1994. Selected papers from an international conference held in Fontainebleu, France, in September 1994.

VanDroogenbroeck:1994:IMO

[Van94a] M. Van Droogenbroeck. On the implementation of morphological operations. In Serra and Soille [SS94], pages 241– 248. ISBN 0-7923-3093-5. LCCN TA1637 .M37 1994. URL http:/

/www.ulg.ac.be/telecom/publi/
publications/mvd/ISMM94-1.pdf;
 http://www.ulg.ac.be/telecom/
publi/publications/mvd/ISMM94-

1.zip. Selected papers from an international conference held in Fontainebleu, France, in September 1994.

VanDroogenbroeck:1994:PEB

[Van94b] M. Van Droogenbroeck. Polymorphism, an extension of binary opening. In Serra and Soille [SS94], pages 1–2. ISBN 0-7923-3093-5. LCCN TA1637 .M37 1994. URL http:/

                                        /www.ulg.ac.be/telecom/publi/
                                        publications/mvd/ISMM94-2.pdf;
                                         http://www.ulg.ac.be/telecom/
                                        publi/publications/mvd/ISMM94-

2.zip. Selected papers from an international conference held in Fontainebleu, France, in September 1994.

VanDroogenbroeck:1994:TIN

[Van94c] M. Van Droogenbroeck. Traitement d’images numériques au moyen d’algorithmes utilisant la morphologie mathématique et la notion d’objet : application au codage. PhD thesis, Catholic University of Louvain, Louvain, Belgium, May 1994. URL

                                        http://www.ulg.ac.be/telecom/
                                        publi/publications/mvd/these/
                                        ;   http://www.ulg.ac.be/telecom/
                                        publi/publications/mvd/these/
                                        theseMVD.pdf;   http://www.ulg.
                                        ac.be/telecom/publi/publications/
                                        mvd/these/theseMVD2.zip.

VanDroogenbroeck:1995:MIS

[Van95] M. Van Droogenbroeck. A methodology for image segmentation based on the spectral content of regions. In Workshop on Nonlinear Signal and Image Processing, pages 266–269. IEEE, Neos Marmaras, June 1995. URL

                                        http://www.ulg.ac.be/telecom/
                                        publi/publications/mvd/ieee_
                                        1995.pdf;   http://www.ulg.ac.

be/telecom/publi/publications/
mvd/ieee_1995.zip.

VanDroogenbroeck:2001:ATI

[Van01a] M. Van Droogenbroeck. Acquisition et traitement de l’image. Notes de cours, Université de Liège, avril 2001. 225 pp. URL

http://www.ulg.ac.be/telecom/
teaching/notes/totaliBIO/index.

html. PDF available on request.

VanDroogenbroeck:2001:ACS

[Van01b] M. Van Droogenbroeck. Analyse et conception des systèmes de télécommunications. Notes de cours, Université de Liège, mars 2001.

222 pp. URL http://www.ulg.
ac.be/telecom/teaching/notes/
total2/index.html. PDF available

on request.

VanDroogenbroeck:2001:TNI

[Van01c] M. Van Droogenbroeck. Traitement numérique des images. Notes de cours, Université de Liège, septembre 2001. 225 pp. URL

http://www.ulg.ac.be/telecom/
teaching/notes/totali/index.

html. PDF available on request.

VanDroogenbroeck:2002:AOB

[Van02a] M. Van Droogenbroeck. Algorithms for openings of binary and label images with rectangular structuring elements. In H. Talbot and R. Beare, editors, Mathematical morphology, pages 197–207. CSIRO Publishing, Sydney, Australia, April 2002. URL http:/

/www.ulg.ac.be/telecom/publi/
publications/mvd/ismm2002.pdf.

VanDroogenbroeck:2002:PTA

[Van02b] M. Van Droogenbroeck. Principes des télécommunications analogiques et numériques. Notes de cours, Université de Liège, janvier 2002.

                                        229 pp. URL http://www.ulg.
                                        ac.be/telecom/teaching/notes/
                                        total1/index.html. PDF available

on request.

VanDroogenbroeck:2002:TDM

[Van02c] M. Van Droogenbroeck. Technologies du multimédia, de l’Internet et des télécommunications. Notes de cours, Institut Supérieur de Commerce (ISC) Saint-Louis, mars 2002.

                                        183 pp. URL http://www.ulg.
                                        ac.be/telecom/teaching/notes/
                                        multimedia/index.html.  PDF

available on request.

VanDroogenbroeck:2002:TO

[Van02d] M. Van Droogenbroeck. Télécommunications et ordinateurs. Notes de cours, Université de Liège, janvier 2002.

                                        205 pp. URL http://www.ulg.
                                        ac.be/telecom/teaching/notes/
                                        total0/index.html. PDF available

on request.

VanDroogenbroeck:2005:MEO

[VB05] M. Van Droogenbroeck and M. J.

Buckley. Morphological erosions and openings: fast algorithms based on anchors. Journal of Mathematical Imaging and Vision: JMIV, 22(2–3):121–142, 2005. CODEN JMIVEK. ISSN 0924-9907 (print), 1573-7683 (electronic).

VanDroogenbroeck:2002:EBT

[VD02] M. Van Droogenbroeck and J. Delvaux. An entropy based technique for information embedding in images. In IEEE Benelux Signal Processing Symposium, Leuven, Belgium, pages 81–84. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, March 2002. URL http:/

/www.ulg.ac.be/telecom/publi/
publications/mvd/sps-034/index.
html;   http://www.ulg.ac.be/
telecom/publi/publications/mvd/
sps-034.pdf.

VanDroogenbroeck:1996:FCM

[VT96] M. Van Droogenbroeck and H. Talbot. Fast computation of morphological operations with arbitrary structuring elements. Pattern Recognition Letters, 17(14):1451– 1460, 1996. CODEN PRLEDG. ISSN 0167-8655 (print), 1872-7344 (electronic). URL http:/

/www.ulg.ac.be/telecom/publi/
publications/mvd/anydila.pdf.

VanDroogenbroeck:1992:ISD

[VV92] M. Van Droogenbroeck and L. Vandendorpe. Integration of the selective deconvolution algorithm into a multiresolution scheme. In James R. Sullivan, Majid Rabbani, and Benjamin M. Dawson, editors, Image processing algorithms and techniques III: 10–13 February 1992, San Jose, California, volume 1654 of SPIE proceedings series, pages 225–235. SPIE Optical Engineering Press, Bellingham, WA, USA, February 1992. ISBN 0-8194-0811-5. LCCN TA1632 .I387 1992; TS510 .P632 v. 1657.

VanTomme:1996:NVB

[VVD96] E. Van Tomme, M. Van Droogenbroeck, and P. Debie. Narrowband versus broadband for tele-education: a critical evaluation. In International Conference on Technology Supported Learning, pages 200–207. Online Educa Berlin, Berlin, Germany, November 1996.